OpenStudy (anonymous):

Find the right end point for f(x)=-2-cos(2x-pi).

OpenStudy (anonymous):

Could also be written as -cos(2x-pi)-2

zepdrix (zepdrix):

So our starting point will be when our angle is zero.\[\Large\rm 2x-\pi=0\]Solving for x gives us:\[\Large\rm x=\frac{\pi}{2}\]That's probably what you got for your starting point, yes? Our ending point will be when we've completed one revolution, in other words, when our angle is 2pi,\[\Large\rm 2x-\pi=2\pi\]

zepdrix (zepdrix):

Are you listing your start and end points as ordered pairs? Or just the x-coordinates?

OpenStudy (anonymous):

Just like in the other problems. I did pi/2 for my phase shift (starting/left end point).

OpenStudy (anonymous):

Right end point = c/b(phase shift) + 2pi/b(period).

zepdrix (zepdrix):

Mmmm ok yes, if we're going to stick with that plan, yes yes yes :) Looks good.

zepdrix (zepdrix):

\[\Large\rm \text{right end point}=\frac{\pi}{2}+\frac{2\pi}{2}\]So something like that, yah? :d

OpenStudy (anonymous):

Right. Then I went a step further by getting the sum. But technically what I got was an improper fractions, which isn't unheard of, but have not come across in this situation.

zepdrix (zepdrix):

Oh ok. Yah that's one of those things to get used to in big boy math: `We don't like mixed numbers, those are bad`. Improper fractions are good :)

zepdrix (zepdrix):

You could say that it's `1 and a half` pi's. But we don't like mixed numbers :o no no no! If you're more comfortable with decimals, it's ok to call it 1.5pi

zepdrix (zepdrix):

Because 3 divided by 2 = 1.5

zepdrix (zepdrix):

I guess I'm making an assumption. I'm hoping that you ended up with 3pi/2, and not 3pi/4 :)

OpenStudy (anonymous):

3pi/2 is what I ended up with and I think she doesn't want us using decimals in this case. So, yay, my math was right!

zepdrix (zepdrix):

yay! :)

OpenStudy (anonymous):

:D